OpenStudy (anonymous):

what are the x intercepts : x^3 -9x^2 +27x -27

OpenStudy (anonymous):

Equate x^3 -9x^2 +27x -27 to zero Thus, x^3 -9x^2 +27x -27 = 0 Solve for x. (-x + 3)(-x^2 + 6x - 9) = 0 (-x + 3)(-1)(x^2 - 6x + 9) = 0 (3 - x)(x - 3)^2 = 0 x1 = x2 = x3 = 3 x = 3 has a multiplicity of 3